# AUDETER: A Large-scale Dataset for Deepfake Audio Detection in Open Worlds

### v1

@@ Abstract.
Speech generation systems can produce remarkably realistic vocalisations that are often indistinguishable from human speech, posing significant authenticity challenges. Although numerous deepfake detection methods have been developed, their effectiveness in real-world environments remains unrealiable due to the domain shift between training and test samples arising from diverse human speech and fast evolving speech synthesis systems. This is not adequately addressed by current datasets, which lack real-world application challenges with diverse and up-to-date audios in both real and deepfake categories. To fill this gap, we introduce AUDETER (AUdio DEepfake TEst Range), a large-scale, highly diverse deepfake audio dataset for comprehensive evaluation and robust development of generalised models for deepfake audio detection. It consists of over 4,500 hours of synthetic audio generated by 11 recent TTS models and 10 vocoders with a broad range of TTS/vocoder patterns, totalling 3 million audio clips, making it the largest deepfake audio dataset by scale. Through extensive experiments with AUDETER, we reveal that i) state-of-the-art (SOTA) methods trained on existing datasets struggle to generalise to novel deepfake audio samples and suffer from high false positive rates on unseen human voice, underscoring the need for a comprehensive dataset; and ii) these methods trained on AUDETER achieve highly generalised detection performance and significantly reduce detection error rate by 44.1% to 51.6%, achieving an error rate of only 4.17% on diverse cross-domain samples in the popular In-the-Wild dataset, paving the way for training generalist deepfake audio detectors. AUDETER is available on GitHub.
@@ 1. Introduction
Deepfake audio detection is the task of identifying audio generated by speech synthesis models, such as Text-To-Speech (TTS) systems and vocoders. There has been a long history of developing detection models for fake audio due to the significance of its real-world applications such as authentication in forensics, misinformation detection on social media, and voice biometric security systems.
Numerous detection methods have been developed and demonstrated to be effective when evaluated against current benchmark datasets, such as ASVSpoof (Wang et al., 2020; Yamagishi et al., 2021; Wang et al., 2024a) and In-the-Wild (Müller et al., 2022). However, the problem of open-world detection remains a major challenge, i.e., detecting deepfake audio samples generated by novel speech synthesis systems (Shen et al., 2018) that are not represented in the training samples, together with human voices with different acoustic features and artefacts. This is because most existing methods treat the detection problem as a closed-set binary classification problem and are optimised to fit the training dataset with limited audio patterns. They fail to consider potential novel patterns at inference time, which results in unsatisfactory generalisation to novel audio samples, especially when deployed in open-world detection settings. Figure 1 illustrates this limitation, showing that a state-of-the-art XLR-S-based detector (Tak et al., 2022) fails at open-world detection on two subsets of our dataset, incorrectly mapping test samples to the wrong regions due to poor generalisation to both deepfake audio samples generated by recent speech models and diverse human voices.
@@
Current datasets have become increasingly limited in covering the diverse and up-to-date deepfake speech patterns. While existing models achieve high performance on these datasets, they do not adequately evaluate the realistic challenges for the detection in open worlds, as previously mentioned. To improve and promote evaluation for open-world detection, we introduce AUDETER (AUdio DEepfake TEst Range), a large-scale deepfake audio detection dataset that collects audio samples generated by a wide variety of speech synthesis models and multiple sources of human voices. Table 1 compares AUDETER with the existing datasets. AUDETER contains 4,682 hours of deepfake audio generated using 21 speech synthesis systems, including 10 recent Text-to-Speech (TTS) models, corresponding to 4 human voice corpora, making it much larger and more diverse than previous datasets. A key advantage is that for each real audio sample, we provide corresponding fake audio generated by all synthesis systems using matching scripts, allowing for systematic, balanced evaluation with consistent structure. By combining different AUDETER components, we can simulate various domain shifts in open-world detection.
@@
Dataset # Audio Clips Total Hours Diverse Real # Synthetic Models # TTS Models # Identical Script ASV 2019 (Wang et al., 2020) 312K 60-70 N 19 N N ASV 2021 DF (Yamagishi et al., 2021) 612K 100-120 N 13 N N In-the-Wild (Müller et al., 2022) 31.8K 17.2 N - N N WaveFake (Frank and Schönherr, 2021) 137K 175 N 6 N Y LibriSeVoc (Sun et al., 2023) - 126.41 N 6 N Y AUDETER (ours) 3M 4,681.9 25 21 11 Y
Through extensive evaluation using AUDETER, we find that existing detection models trained on current datasets experience significant performance drops when used for open-world detection, further demonstrating the limitations in current benchmarks and training resources. Our key finding is that while generalising to all possible acoustic patterns is infeasible, speech synthesis systems share similar audio patterns. By combining a set of diverse, representative systems, we can achieve reasonably universal generalisation to most systems. In addition, training with large-scale data incorporating diverse sources of human voices will improve the domain transferability of learned deepfake patterns against domain shift in real audio. AUDETER’s consistent structure and systematic design enables understanding of the relationships between systems and human voice sources for developing more effective training data compositions.
Owing to the high diversity and large scale, AUDETER serves as a valuable resource for training open-world detectors and enables a data-centric approach to improve detection performance. All models trained on AUDETER achieved significantly improved performance compared to their pretrained versions, achieving an Equal Error Rate of 4.17% using our XLR-SLS detector. As large audio backbones become increasingly popular, the scale of our dataset can support training these data-demanding models containing millions of parameters. Our main contributions are as follows:
We identify and analyse fundamental limitations of existing deepfake audio detection methods in open-world scenarios, demonstrating through systematic evaluation that these methods are essentially closed-set binary classification approaches, failing to generalise to novel speech synthesis systems and diverse human voice characteristics not represented in training data.
@@
We introduce AUDETER, a large-scale deepfake audio detection dataset comprising 4,682 hours of synthetic audio generated by 21 recent speech synthesis systems across 4 human voice corpora, which is substantially larger and more diverse than existing benchmarks, with systematic design enabling comprehensive evaluation of domain shifts.
We train three popular detector architectures using AUDETER and show that AUDETER effectively provides diverse audio samples as a data-centric approach for improving open-world detection, achieving superior performance compared to SOTA methods.
@@ 2. Related Work and Preliminaries
2.1. Synthetic Audio Detection
@@
Detection models for deepfake audio have been extensively studied (Wang et al., 2020; Yamagishi et al., 2021; Frank and Schönherr, 2021; Yi et al., 2022; Müller et al., 2022; Sun et al., 2023; Wang et al., 2024a; Zhang et al., 2025; Ren et al., 2025). Earlier detection frameworks concentrated on exploring different types of features, such as short-term spectral (Xiao et al., 2015; Tian et al., 2016), long-term spectral (Alegre et al., 2013b, a; Sahidullah et al., 2015), prosodic (Pal et al., 2018; Xue et al., 2022; Wang et al., 2023), and deep features (Yu et al., 2017; Sailor et al., 2017; Cheuk et al., 2020; Fu et al., 2022; Zeghidour et al., 2018; Ravanelli and Bengio, 2018). Another direction leverages end-to-end deep neural networks for enhanced performance with a diverse range of detector architectures (Wu et al., 2020; Liu et al., 2023; Tak et al., 2021a; Jung et al., 2022; Ge et al., 2021b). More recent works employ pretrained audio models such as Wav2Vec2.0 (Baevski et al., 2020) as backbones, leveraging their substantially stronger capacity with millions of parameters and extensive pretraining knowledge in the audio domain for feature extraction, These backbones (Babu et al., 2021; Conneau et al., 2020) can then be combined in a pipeline with deep neural networks similar to the previously mentioned architectures serving as scoring networks. They have demonstrated significant performance improvements and represent the current state-of-the-art (Yang et al., 2024; Tak et al., 2022; Zhang et al., 2024). However, despite their enhanced learning capacity, they rely on supervised learning frameworks, and therefore are not readily adaptable to novel audio patterns.
2.2. Speech Synthesis Models
2.2.1. End-to-end TTS Systems
@@
TTS systems convert text input to audio waveforms through neural networks. Earlier approaches like Tacotron (Wang et al., 2017), WaveNet (Oord et al., 2016), and FastSpeech (Ren et al., 2019) follow a two-stage workflow, in which text inputs are first converted to acoustic features, which are then processed by vocoders to generate waveforms. More sophisticated generative models are used to improve this process such as flow-based generation (Glow-TTS) (Kim et al., 2020) and end-to-end variational approaches (VITS) (Kim et al., 2021). More recent models such as YourTTS (Casanova et al., 2022) and OpenVoice (Qin et al., 2023) enable voice cloning with basic emotion and style control. However, these methods rely on phonetic alignments for generation and suffer from limited emotional realism and poor prosodic modelling capabilities such as intonation. More recent TTS systems (Du et al., 2024; Wang et al., 2025; Liao et al., 2024; Suno AI, 2023; Kim et al., 2021) leverage large language models and large-scale pretraining and significantly improved the quality and realism of speech generation.
2.2.2. Vocoders
@@
Vocoders convert intermediate acoustic features (such as mel-spectrograms) into audio waveforms. Traditional methods (Morise et al., 2016; Degottex et al., 2011) used signal processing techniques, but neural vocoders have substantially improved audio quality. Early neural approaches (Oord et al., 2016; Mehri et al., 2016) use autoregressive generation but suffer from slow inference. Subsequent developments like (Prenger et al., 2019) and Parallel WaveGAN (Yamamoto et al., 2020) introduced parallel generation for faster synthesis. GAN-based vocoders such as MelGAN(Kumar et al., 2019), HiFi-GAN (Kong et al., 2020), and UnivNet (Jang et al., 2021) further improved efficiency and quality through adversarial training. More recent vocoders like BigVGAN (Lee et al., 2022) and Vocos (Siuzdak, 2023) have incorporated advanced architectures and training techniques to achieve state-of-the-art audio fidelity. Modern vocoders can generate audio that is nearly indistinguishable from real human speech.
@@ 2.2.3. Deepfake Audio Datasets
ASVSpoof series datasets (Wang et al., 2020; Yamagishi et al., 2021) are widely used for deepfake audio detection. ASVSpoof 2019 focuses on text-to-speech and voice conversion attacks, while ASVSpoof 2021 expanded to include more diverse spoofing methods.In-the-Wild (Müller et al., 2022) is a common choice for evaluating cross-domain detection, but its small scale and unspecified generation methods make it unsuitable for training and fine-grained evaluation. WaveFake (Frank and Schönherr, 2021) collected synthetic audio from various vocoders models, and LibriSeVoc contributed additional evaluation data by synthesising speech from LibriSpeech (Sun et al., 2023) using various vocoding techniques. However, they are relatively small-scale and primarily feature vocoders and legacy speech generation methods, with limited representation of recent end-to-end TTS systems.
@@ 2.3. Problem Statement
The objective of deepfake audio detection is to train a scoring mechanism that assigns scores to audio samples that are indicative of their authenticity. Within a normalised scale of 0 to 1, where 0 and 1 are arbitrarily set to represent fake and real, the learning scoring function should ideally satisfy , for any real audio clip and . In practice, a threshold is applied to the score as a decision boundary, classifying samples as real if and fake otherwise. The detection performance is commonly evaluated using the Equal Error Rate (EER), which is defined as the error rate at the decision threshold where the false acceptance rate (FAR) equals the false rejection rate (FRR):
where is the threshold at which the two rates are equal. A lower EER indicates better overall detection performance.
Most existing methods implement using DNNs. Open-world deepfake detection refers to detecting deepfake audio at test time when facing synthetic audio generated by novel speech synthesis systems not seen during training, as well as human voices with different styles and characteristics.
@@ 3.1. Design Motivations
The AUDETER dataset is motivated to address key limitations in current deepfake detection evaluation and bridge the gap of large-scale datasets for training generalised detection models for open-world detection.
3.1.1. Enhanced Evaluation for Open-world Detection.
Existing datasets lack comprehensive coverage of recent synthesis methods, particularly pretrained TTS models and recent vocoders that have significantly advanced quality and realism. Second, their human speech audios are limited in number and diversity. As shown in Section 4.2, current models suffer from significant performance degradation under human voice shift and novel speech synthesis systems. Therefore, we generate deepfake audio using 21 recent TTS and vocoder models to address this gap, corresponding to human speech from four corpora with different style.
@@
The performance of existing models deteriorates under domain shift because they are fitted to limited training audio samples. A data-centric approach toward building generalised detection models involves training with diverse real and deepfake audio samples. This requires strong model learning capacity and large amounts of training data. AUDETER offers diverse high-quality data that can be used for large-scale training and potential self-supervised learning.
@@ 3.2. Dataset Overview
Collection Subset Partition Patterns # Audio / Patt. Total Hrs TTS In-the-Wild Bona-fide 15 19,784 311.5 Common Voice Val 15 16,372 275.0 Test 15 16,372 265.3 People Speech Val 15 18,622 493.5 Test 15 34,898 909.4 MLS Dev 15 3,807 212.7 Test 15 3,769 209.1 Vocoder In-the-Wild Bona-fide 10 19,784 207.6 Common Voice Val 10 16,372 266.7 Test 10 16,372 264.8 People Speech Val 10 18,622 331.7 Test 10 34,898 598.1 MLS Dev 10 3,807 156.7 Test 10 3,769 154.9
TTS Systems CosyVoice (2025) (Du et al., 2024), Zonos (2025) (Zyphra-Zonos, 2024), SparkTTS (2025) (Wang et al., 2025), F5-TTS (2025) (Chen et al., 2024), Fish-Speech (2024) (Liao et al., 2024), OpenVoice V2 (2023) (Qin et al., 2023), ChatTTS (2024) (2noise, 2024), XTTS v2 (2024) (Casanova et al., 2024), Bark (2023) (Suno AI, 2023), YourTTS (2022) (Casanova et al., 2022), VITS (2021) (Kim et al., 2021) Vocoders BigVGan (2022) (Lee et al., 2022), BigVSan (2024) (Shibuya et al., 2024), Vocos (Siuzdak, 2023) (2023), UnivNet (2021) (Jang et al., 2021), HiFi-GAN (2020) (Kong et al., 2020):, MelGAN (2019) (Kumar et al., 2019) , Full-band MelGAN (kan bayashi, 2019), Multi-band MelGAN (Yang et al., 2021), Parallel WaveGAN (Yamamoto et al., 2020), Style MelGAN (Mustafa et al., 2021).
We summarise the structure of AUDETER in Table 2. AUDETER consists of two collections: the TTS and the Vocoder collection, where the TTS collection contains multiple versions of deepfake audio generated using recent end-to-end TTS systems that pronounce the identical scripts as their corresponding real audio. The Vocoder collection contains vocoded audio of the real audio. Each collection is further divided into four subsets according to the source of real audio. For instance, the validation subset of the Common Voice partition consists of 16,372 real audio samples and 16 versions × 16,372 TTS-generated audio samples and 10 versions × 16,372 vocoded audio samples.
3.2.1. The Sources of Real Speech.
We include real audio samples from 4 corpora to create diverse human speech distributions. Specifically, we include all English real audio samples from the In-the-wild dataset, validation and test partitions of the Common Voice and People’s Speech Dataset, and dev and test partitions from the multilingual LibriSpeech dataset, which provides complementary characteristics. In-the-Wild offers varied recording conditions and speakers, Common Voice provides crowdsourced read speech with diverse accents, People’s Speech contains broadcast-quality professional recordings, and multilingual LibriSpeech contributes clean audiobook recordings across multiple languages, covering comprehensive human speech variability.
@@
The speech models used for synthetic audio generation are summarised in Table 3. We employ 11 popular open-source TTS systems for text-to-waveform generation. For OpenVoice V2, we use 5 of its default speakers to generate 5 versions of synthetic audio for studying the effect of voice reference. We also employ 10 recent vocoders for vocoded audio. The models are selected based on their popularity and recency.

### v2

@@ Abstract.
Speech synthesis systems can now produce highly realistic vocalisations that pose significant authenticity challenges. Despite substantial progress in deepfake detection models, their real-world effectiveness is often undermined by evolving distribution shifts between training and test data, driven by the complexity of human speech and the rapid evolution of synthesis systems. Existing datasets suffer from limited real speech diversity, insufficient coverage of recent synthesis systems, and heterogeneous mixtures of deepfake sources, which hinder systematic evaluation and open-world model training. To address these issues, we introduce AUDETER (AUdio DEepfake TEst Range), a large-scale and highly diverse deepfake audio dataset comprising over 4,500 hours of synthetic audio generated by 11 recent TTS models and 10 vocoders, totalling 3 million clips. We further observe that most existing detectors default to binary supervised training, which can induce negative transfer across synthesis sources when the training data contains highly diverse deepfake patterns, impacting overall generalisation. As a complementary contribution, we propose an effective curriculum-learning-based approach to mitigate this effect. Extensive experiments show that existing detection models struggle to generalise to novel deepfakes and human speech in AUDETER, whereas XLR-based detectors trained on AUDETER achieve strong cross-domain performance across multiple benchmarks, achieving an EER of 1.87% on In-the-Wild. AUDETER is available on GitHub.
@@ 1. Introduction
Deepfake audio detection (DFAD) is the task of identifying audio generated by speech synthesis models, such as Text-To-Speech (TTS) systems and vocoders. There has been a long history of developing DFAD models due to the significance of its real-world applications such as authentication in forensics, misinformation detection on social media, and voice biometric security systems.
State-of-the-art (SOTA) DFAD models have shown strong performance against current benchmark datasets, such as ASVSpoof (Wang et al., 2020; Yamagishi et al., 2021; Wang et al., 2024b) and In-the-Wild (ITW) (Müller et al., 2022). However, their performance in real-world scenarios remains unreliable, particularly when detecting deepfake audio generated by novel speech synthesis systems not seen during training, as well as human voices with different acoustic characteristics and artefacts. This is because most DFAD models are trained under a closed-world binary classification formulation, which fits only the limited acoustic and synthetic patterns present in the training data and consequently degrades open-world generalisation, as illustrated in Fig. 1. We refer to open-world detection as the setting where both real and synthetic audio samples at inference time may originate from previously unseen distributions or synthesis systems not explicitly included in training.
@@
Current datasets are becoming increasingly limited in covering diverse and recent deepfake speech patterns. As a result, they fail to adequately challenge modern DFAD models for open-world detection. To improve open-world DFAD evaluation and model training, we introduce AUDETER (AUdio DEepfake TEst Range), a large-scale deepfake audio detection dataset that consists of audio samples systematically created using a wide variety of speech synthesis models and multiple sources of human voices. Table 1 compares AUDETER with the existing datasets. AUDETER contains 4,682 hours of deepfake audio generated using 21 speech synthesis systems, including 10 recent TTS models, corresponding to 4 human voice corpora, making it substantially larger and more diverse than previous datasets. In particular, each real audio sample is paired with matched synthetic versions from all 21 synthesis systems, enabling balanced evaluation and controlled analysis of domain shifts in open-world detection. Moreover, the quality of all generated audio is rigorously assessed for intelligibility and naturalness (see Fig. 8 for AUDETER’s audio generation pipeline).
@@
Dataset # Audio Clips Total Hours Diverse Real # Synthetic Models # TTS Models # Matching Script ASV 2019 (Wang et al., 2020) 312K 60-70 N 19 N N ASV 2021 DF (Yamagishi et al., 2021) 612K 100-120 N 13 N N In-the-Wild (Müller et al., 2022) 31.8K 17.2 N - N N WaveFake (Frank and Schönherr, 2021) 137K 175 N 6 N Y LibriSeVoc (Sun et al., 2023) - 126.41 N 6 N Y DFADD (Du et al., 2024a) 1K 221 N 5 Y N SONAR (Li et al., 2025) 2K 8 N 8 Y N AUDETER (ours) 3M 4,681.9 4 21 11 Y
AUDETER not only serves as a challenging and systematic evaluation benchmark, but a valuable resource that offers diverse coverage at scale for training DFAD models. For evaluation, our extensive experiments on AUDETER reveal significant performance degradation of models trained on existing datasets that include recent, novel systems, underscoring the limitations of current benchmarks and training data. AUDETER’s consistent and uniform structure enables systematic cross-system analysis, providing a flexible foundation for future investigations into DFAD generalisation. We present representative findings in Section 4.3.
For training, detection models trained on AUDETER consistently outperform their pretrained counterparts under the same training setup, as well as leading baseline detectors, in both in-domain and cross-dataset evaluations. This highlights AUDETER’s comprehensive and representative data coverage and its effectiveness as a data-centric approach to improving detection performance. In addition, AUDETER-trained models exhibit increased robustness to real-world challenges such as background noise. Specifically, using a SOTA DFAD model architecture, XLR-SLS (Zhang et al., 2024), we retrain the detector on AUDETER and achieve an Equal Error Rate (EER) of 1.87% on the ITW dataset, demonstrating that AUDETER provides sufficient training coverage for data-demanding backbone models.
As a complementary contribution, we identify a key limitation of the default closed-world supervised training approach commonly used by existing detection models, which treats all deepfake audio as a single homogeneous class without explicitly modelling differences across speech synthesis systems. Enabled by AUDETER’s diverse synthesis patterns, we find that directly training under this formulation on highly diverse speech models can induce negative transfer, degrading cross-domain performance. To further study this effect, we extend cross-domain evaluation to two additional recent datasets, revealing that systems with strong synthesis fingerprints can disproportionately harm overall cross-domain generalisation. To alleviate this issue, we cast learning from diverse deepfake data as a curriculum learning problem implemented via a two stage pipeline. In stage one, we identify systems with strong synthesis fingerprints that often dominate the training signal using a linear probing test and train a base detector on the remaining data to learn universal, system invariant representations. In stage two, the model is adapted to all speech systems, with the stage one backbone used as a teacher to regularise adaptation and preserve general cues that cannot be enforced through data labels. This leads to balanced and substantially improved cross-domain performance across multiple benchmarks. Our main contributions are as follows:
@@
We introduce AUDETER, a large-scale deepfake audio detection dataset comprising 4,682 hours of synthetic audio generated by us using 21 recent speech synthesis systems across four human voice corpora. AUDETER provides substantially greater scale and diversity than existing datasets, supporting improved evaluation and model training for open-world detection.
Through extensive evaluation, AUDETER is shown to effectively challenge existing models, exposing their limited suitability for open-world detection. As a training resource, AUDETER enables significant improvements over prior SOTA performance.
We identify a key limitation of closed-world training where dominant synthesis fingerprints hinder generalisation, and address this with a curriculum learning based method that significantly improves cross-domain performance.
@@ 2. Related Work and Preliminaries
2.1. Related Work
2.1.1. Deepfake Audio Detection.
@@ 2.1. Related Work
DFAD models have been extensively studied (Wang et al., 2020; Yamagishi et al., 2021; Frank and Schönherr, 2021; Yi et al., 2022; Müller et al., 2022; Sun et al., 2023; Wang et al., 2024b; Zhang et al., 2025; Ren et al., 2025; Liang et al., 2026). Earlier DFAD models explore various types of features, such as short-term spectral (Xiao et al., 2015; Tian et al., 2016), long-term spectral (Alegre et al., 2013b, a; Sahidullah et al., 2015), prosodic (Pal et al., 2018; Xue et al., 2022; Wang et al., 2023), and deep features (Yu et al., 2017; Sailor et al., 2017; Cheuk et al., 2020; Fu et al., 2022; Zeghidour et al., 2018; Ravanelli and Bengio, 2018). A wide range of deep neural networks (DNN) architectures (Wu et al., 2020; Liu et al., 2023; Tak et al., 2021b; Jung et al., 2022; Ge et al., 2021b) have also been studied for enhanced feature extraction. More recent works employ pretrained audio models such as Wav2Vec 2.0 (Baevski et al., 2020) as backbones, leveraging their greater capacity and extensive audio-domain pretraining for improved representation learning. These backbones (Babu et al., 2021; Conneau et al., 2020) are commonly paired with DNNs as scoring models and represent the current state-of-the-art (Yang et al., 2024; Tak et al., 2022; Zhang et al., 2024). However, these models are trained as closed-set classifiers and often fail to generalise to novel samples at inference time.
2.1.2. Speech Synthesis Models
@@
End-to-end TTS Systems. TTS systems convert text input to audio waveforms through neural networks. Earlier approaches like Tacotron (Wang et al., 2017), WaveNet (Oord et al., 2016), and FastSpeech (Ren et al., 2019) follow a two-stage workflow, in which text inputs are first converted to acoustic features, which are then processed by vocoders to generate waveforms. More sophisticated generative models are used to improve this process such as flow-based generation (Glow-TTS) (Kim et al., 2020) and end-to-end variational approaches (VITS) (Kim et al., 2021). More recent models such as YourTTS (Casanova et al., 2022) and OpenVoice (Qin et al., 2023) enable voice cloning with basic emotion and style control. However, these methods rely on phonetic alignments for generation and suffer from limited emotional realism and poor prosodic modelling capabilities such as intonation. More recent TTS systems (Du et al., 2024b; Wang et al., 2025; Liao et al., 2024; Suno AI, 2023; Kim et al., 2021) leverage large language models and large-scale pretraining, significantly improving speech quality and realism.
@@
Vocoders. Vocoders convert intermediate acoustic features (such as mel-spectrograms) into audio waveforms. Traditional methods (Morise et al., 2016; Degottex et al., 2011) used signal processing techniques, but neural vocoders have substantially improved audio quality. Early neural approaches (Oord et al., 2016; Mehri et al., 2016) use autoregressive generation but suffer from slow inference. Subsequent developments like (Prenger et al., 2019) and Parallel WaveGAN (Yamamoto et al., 2020) introduced parallel generation for faster synthesis. GAN-based vocoders such as MelGAN (Kumar et al., 2019), HiFi-GAN (Kong et al., 2020), and UnivNet (Jang et al., 2021) further improved efficiency and quality through adversarial training. More recent vocoders like BigVGAN (Lee et al., 2022) and Vocos (Siuzdak, 2023) have incorporated advanced architectures and training techniques to achieve state-of-the-art audio fidelity. Modern vocoders can generate audio that is nearly indistinguishable from real human speech.
@@ 2.1.3. Deepfake Audio Datasets
ASVSpoof series datasets (Wang et al., 2020; Yamagishi et al., 2021) are widely used for deepfake audio detection. ASVSpoof 2019 focuses on TTS and voice conversion attacks, while ASVSpoof 2021 expanded to include more diverse spoofing methods.In-the-Wild (Müller et al., 2022) is a common choice for evaluating cross-domain detection, but its small scale and unspecified generation methods make it unsuitable for training and fine-grained evaluation. WaveFake (Frank and Schönherr, 2021) and LibriSeVoc focus mainly on vocoder-based synthesis and remain limited in scale. DFADD (Du et al., 2024a) and SONAR (Li et al., 2025) include multiple TTS systems but are primarily evaluation-focused and not designed for large-scale training.
@@ 2.2. Problem Statement
Deepfake audio detection aims to learn a scoring function that assigns authenticity scores to audio samples, such that . A decision threshold is applied to classify samples as real if and fake otherwise. Detection performance is commonly evaluated using the Equal Error Rate (EER), defined as the error rate at the decision threshold where the false acceptance rate (FAR) equals the false rejection rate (FRR). Lower EER indicates better detection performance. Most existing methods implement using deep neural networks learned through supervised training, while open-world deepfake detection requires generalisation to unseen synthesis systems and diverse human speech styles.
@@ 3.1. Design Motivations
AUDETER is designed to address key limitations in existing deepfake audio benchmarks, both for evaluating open-world detection and for supporting the training of generalised detectors at scale.
3.1.1. Enhanced Open-world Evaluation
Existing datasets provide limited coverage of recent speech synthesis methods and lack sufficient diversity in real speech. As shown in Section 4.2, this leads to significant performance degradation under novel synthesis systems and human voice domain shifts. To address this, AUDETER includes synthetic audio from 21 recent TTS and vocoder models, paired with real speech from four large-scale corpora covering diverse speaking styles and conditions. Consistent with our open-world definition, AUDETER is structured to allow controlled evaluation under both unseen synthesis systems and unseen real-speech distributions.
@@ 3.1.1. Enhanced Open-world Evaluation
Existing models struggle under domain shifts due to limited training data. A data-centric approach to generalised detection requires large-scale training data capturing diverse real and deepfake patterns, which is fundamental for effective model learning and is provided by AUDETER. Beyond scale and diversity, AUDETER’s balanced and structured design enables systematic study of training strategies for open-world generalisation under controlled distribution shifts.
@@ 3.2. Dataset Overview
Collection Subset Partition Patterns # Audio / Patt. Total Hrs TTS Celebrity (CE) Test 15 19,784 311.5 CrowdSource (CS) Train 15 16,372 275.0 Test 15 16,372 265.3 US Congress (UC) Train 15 18,622 493.5 Test 15 34,898 909.4 Audiobook (AB) Train 15 3,807 212.7 Test 15 3,769 209.1 Vocoder Celebrity (CE) Test 10 19,784 207.6 CrowdSource (CS) Train 10 16,372 266.7 Test 10 16,372 264.8 US Congress (UC) Train 10 18,622 331.7 Test 10 34,898 598.1 Audiobook (AB) Train 10 3,807 156.7 Test 10 3,769 154.9
TTS Syst. CosyVoice (2025) (Du et al., 2024b), Zonos (2025) (Zyphra-Zonos, 2024), SparkTTS (2025) (Wang et al., 2025), F5-TTS (2025) (Chen et al., 2024), Fish-Speech (2024) (Liao et al., 2024), OpenVoice V2 (2023) (Qin et al., 2023), ChatTTS (2024) (2noise, 2024), XTTS v2 (2024) (Casanova et al., 2024), Bark (2023) (Suno AI, 2023), YourTTS (2022) (Casanova et al., 2022), VITS (2021) (Kim et al., 2021) Vocoders BigVGan (2022) (Lee et al., 2022), BigVSan (2024) (Shibuya et al., 2024), Vocos (Siuzdak, 2023) (2023), UnivNet (2021) (Jang et al., 2021), HiFi-GAN (2020) (Kong et al., 2020):, MelGAN (2019) (Kumar et al., 2019) , Full-band MelGAN (kan-bayashi, 2019), Multi-band MelGAN (Yang et al., 2021), Parallel WaveGAN (Yamamoto et al., 2020), Style MelGAN (Mustafa et al., 2021).
Table 2 summarises the structure of AUDETER, which comprises two collections: a TTS collection and a Vocoder collection. The TTS collection contains synthetic audio generated by recent end to end TTS systems, with each sample aligned to the script of its corresponding real audio, while the Vocoder collection contains vocoded versions of the same real speech. Both collections are partitioned into four subsets based on the real audio source. For example, the validation subset of the UC subset includes 16,372 real samples and 25 corresponding synthetic variants per sample generated via TTS and vocoding. More details on the usage are discussed in Sec. A.4.
3.2.1. Sources of Real Speech.
We include four real speech corpora to capture a range of speaking styles and recording conditions, and name their corresponding AUDETER subsets based on their defining characteristics and provenance, namely Celebrity, CrowdSource, US Congress and Audiobook. The Celebrity, CrowdSource, US Congress, and Audiobook subsets respectively represent in-the-wild public figure speech, user-contributed scripted speech with diverse accents, formal US-accented congressional recordings, and clean read-book speech. We detail the source corpora in Sec. A.3
@@
Table 3 summarises the speech models used for synthetic audio generation, including 11 open source TTS systems and 10 vocoders, selected for their recency and popularity. For OpenVoice V2, five default speakers are used to produce five synthetic variants to study the effect of voice reference.